# Local Dimension Spectrum for Dominated Irreducible Planar Self-Affine Sets

### v1

@@ 2.2 Lyapunov Dimension
We shall consider maps $\theta\colon\Sigma_{*}\to(0,\infty)$ which we refer to as potentials. We say that a potential $\theta$ is sub-multiplicative if $\theta(\mathtt{i}\mathtt{j})\leqslant\theta(\mathtt{i})\theta(\mathtt{j})$ for all $\mathtt{i},\mathtt{j}\in\Sigma_{*}$. A potential $\theta$ is super-multiplicative if the inverse $1/\theta$ is sub-multiplicative. We furthermore say that a potential $\theta$ is almost-multiplicative if there is a constant $C\geqslant 1$ such that $C\theta$ is sub-multiplicative and $C^{-1}\theta$ is super-multiplicative, and multiplicative if the constant $C$ can be chosen to $1$. Let $\mathcal{M}_{\sigma}(\Sigma)$ denote the collection of all $\sigma$-invariant Borel probability measures on $\Sigma$. For a sub-multiplicative potential $\theta$ and $\nu\in\mathcal{M}_{\sigma}(\Sigma)$, we define
@@
The following lemma guarantees that $\Lambda$ is well-defined.
@@
for $\nu$-almost all $\mathtt{i}\in\Sigma$. Furthermore, $\nu\mapsto\Lambda(\theta,\nu)$ defined on $\mathcal{M}_{\sigma}(\Sigma)$ is upper semi-continuous in the $*{}^{*}$ topology.
@@
The second claim follows from Kingman’s ergodic theorem. The third claim is a direct consequence of the first claim as each $\nu\mapsto\frac{1}{n}\sum_{\mathtt{i}\in\Sigma_{n}}\nu([\mathtt{i}])\log\theta(\mathtt{i})$ is continuous. ∎
[…]
@@
We say that a fully supported measure $\mu\in\mathcal{M}(\Sigma)$ is sub-multiplicative if the potential $\mathtt{i}\mapsto\mu([\mathtt{i}])$ is sub-multiplicative. The other definitions on potentials can be used with measures in a similar manner. Regardless, almost-multiplicative measures are more commonly known as quasi-Bernoulli measures and multiplicative measures as Bernoulli measures. The cross-entropy of a sub-multiplicative measure $\mu$ relative to $\nu\in\mathcal{M}_{\sigma}(\Sigma)$ is defined to be
@@
Since $-x\log x\geqslant 0$ for all $x\in(0,1]$, we have $h(\nu)\geqslant 0$. To show the other inequality, fix $n\in\mathbb{N}$ and, as $\log x\leqslant x-1$ for all $x>0$, note that
@@ 2.3 Domination
We say that $\mathsf{A}=(A_{1},\ldots,A_{N})\in GL_{2}(\mathbb{R})^{N}$ is dominated if there exist constants $C>0$ and $10<\tau<$ 1 such that
@@
for all $\mathtt{i}\in\Sigma_{n}$ and $n\in\mathbb{N}$. Note that if $\mathsf{A}$ is dominated, then $\lambda_{2}(\nu)<\lambda_{1}(\nu)$ for all $\nu\in\mathcal{M}_{\sigma}(\Sigma)$. Let $\mathbb{RP}^{1}$ denote the real projective line, which is the set of straight lines going through the origin in $\mathbb{R}^{2}$ and which we identify with $[0,\pi)$. We call a proper subset $\mathcal{C}\subset\mathbb{RP}^{1}$ a multicone if it is a finite union of closed projective intervals. We say that a multicone $\mathcal{C}\subset\mathbb{RP}^{1}$ is strongly invariant for $\mathsf{A}$ if $A_{i}\mathcal{C}\subset\mathcal{C}^{o}$ for all $i\in\{1,\ldots,N\}$, where $\mathcal{C}^{o}$ is the interior of $\mathcal{C}$. For example, the union of first and third quadrants is strongly invariant for any tuple of positive matrices. Notice that if a multicone $\mathcal{C}\subset\mathbb{RP}^{1}$ is strongly invariant for $\mathsf{A}=(A_{1},\ldots,A_{N})$, then the closure of $\mathbb{RP}^{1}\setminus\mathcal{C}$ is a strongly invariant multicone for $(A_{1}^{-1}\ldots,A_{N}^{-1})$. By [9, Theorem B], $\mathsf{A}$ has strongly invariant multicone if and only if $\mathsf{A}$ is dominated. Therefore, if $\mathsf{A}$ is dominated, then the tuple $(A_{1}^{-1}\ldots,A_{N}^{-1})$ of inverses has a strongly invariant multicone. Also if $\mathsf{A}$ is dominated, then [10, Lemma 2.2] (see also [6, Corollary 2.4]) implies that the potential $\mathtt{i}\mapsto\alpha_{1}(\mathtt{i})$ is almost-multiplicative. Since $|\det(A_{\mathtt{i}})|=\alpha_{1}(\mathtt{i})\alpha_{2}(\mathtt{i})$ for all $\mathtt{i}\in\Sigma_{*}$ and the determinant is multiplicative, we see that also $\mathtt{i}\mapsto\alpha_{2}(\mathtt{i})=\alpha_{1}(\mathtt{i})^{-1}|\det(A_{\mathtt{i}})|$ is almost-multiplicative.
@@
Let $\theta$ be an almost-multiplicative potential and $\nu\in\mathcal{M}_{\sigma}(\Sigma)$. If $\nu_{k}\to\nu$ in the $*{}^{*}$ topology, then
@@
Note that $X_{F}$ is perfect unless it is a singleton. See [1, Lemma 2.3] for alternative characterizations of $X_{F}$. If $\nu\in\mathcal{M}_{\sigma}(\Sigma)$ is ergodic, then the measure $\nu_{F}=\Pi_{*}\nu$ on $X_{F}$ is called the Furstenberg measure with respect to $\nu$. Let $T\colon\Sigma\times\mathbb{RP}^{1}\to\Sigma\times\mathbb{RP}^{1}$ be the skew-product defined by
@@
If $\nu$ is Bernoulli, the measure $\nu\times\nu_{F}$ on $\Sigma\times\mathbb{RP}^{1}$ is $T$-invariant and ergodic, if $\nu$ is only quasi-Bernoulli then $\nu\times\nu_{F}$ may not be invariant, but is equivalent to a $T$-invariant ergodic measure; see [5, Theorem 2.2] and [14, Lemma 3.1], respectively.
@@ 2.4 Equilibrium States
As in Lemma 2.1, the existence of the limit above and the equality are guaranteed by Fekete’s Lemma. By [16, Lemma 2.2], we see that
@@
is called the equilibrium state for $\theta$. If $\theta$ is an almost-multiplicative potential, then, by [16, §3], there exists a unique equilibrium state for $\theta$ which furthermore is a fully supported quasi-Bernoulli measure.
@@
Let $(\theta_{k})_{k\in\mathbb{N}}$ be a sequence of sub-multiplicative potentials and let $\nu_{k}\in\mathcal{M}_{\sigma}(\Sigma)$ be an equilibrium state for $\theta_{k}$ for each $k\in\mathbb{N}$. If there exist a measure $\nu\in\mathcal{M}_{\sigma}(\Sigma)$ and a sub-multiplicative potential $\theta$ such that $\nu_{k}\to\nu$ in the $*{}^{*}$ topology and $\theta_{k}(\mathtt{i})^{1/|\mathtt{i}|}\to\theta(\mathtt{i})^{1/|\mathtt{i}|}$ uniformly in $\Sigma_{*}$ as $k\to\infty$, then
@@
Recall that, by [22, Theorem 8.2] and Lemma 2.1, $\limsup_{k\to\infty}h(\nu_{k})\leqslant h(\nu)$ and $\limsup_{k\to\infty}\Lambda(\theta,\nu_{k})\leqslant\Lambda(\theta,\nu)$. If $\varepsilon>0$, then the uniform convergence of $\theta_{k}$ implies that there exists $k_{0}\in\mathbb{N}$ such that $\log\theta(\mathtt{i})-\varepsilon|\mathtt{i}|\leqslant\log\theta_{k}(\mathtt{i})\leqslant\log\theta(\mathtt{i})+\varepsilon|\mathtt{i}|$ for all $\mathtt{i}\in\Sigma_{*}$ and
@@
for all $\mathtt{i}\in\Sigma_{*}$. Since $\alpha(\mathtt{i})\alpha_{2}(\mathtt{i})^{s-1}=\alpha_{1}(\mathtt{i})^{2-s}|\det(A_{\mathtt{i}})|^{s-1}$, the singular value function $\varphi^{s}$ is sub-multiplicative. Therefore, the pressure $P(\varphi^{s})$ is well-defined for all $s\geqslant 0$. By [17, Lemma 2.1], the function $s\mapsto P(\varphi^{s})$ defined on $[0,\infty)$ is continuous and strictly decreasing, convex on intervals $[0,1]$, $[1,2]$, and $[2,\infty)$, and there exists a unique $s\geqslant 0$ such that $P(\varphi^{s})=0$. This unique $s\geqslant 0$ is called the affinity dimension and it is denoted by $\operatorname{dim_{aff}}(\varphi^{s})$.
@@
By [15, Theorems 2.6 and 4.1], there exists an ergodic equilibrium state $\nu$ for $\varphi^{s}$. If $\mathsf{A}$ is dominated, then $\varphi^{s}$ is almost-multiplicative and $\nu$ is quasi-Bernoulli. It is easy to see that an equilibrium state $\nu$ has maximal possible Lyapunov dimension, $\operatorname{dim_{L}}(\nu)=\max\{\operatorname{dim_{L}}(\eta):\eta\in\mathcal{M}_{\sigma}(\Sigma)\}=\operatorname{dim_{aff}}(\varphi^{s})$.
@@ 3 Differentiability of the Pressure
is well-defined and there exists a unique ergodic equilibrium state for $\psi^{q,s}$ which furthermore is a fully supported quasi-Bernoulli measure.
In this section, we prove that the function $(q,s)\mapsto P(\psi^{q,s})$ is differentiable on ×$\mathbb{R}\times(0,\infty)\setminus\{1,2\}$ and determine its partial derivatives. We begin by collecting some elementary properties of the pressure function.
@@
    The function $(q,s)\mapsto P(\psi^{q,s})$ is continuous on ×$\mathbb{R}\times[0,\infty)$.
@@
    For each $q<1$ the function $s\mapsto P(\psi^{q,s})$ is strictly decreasing with $P(\psi^{q,0})\geqslant 0$ and $\lim_{s\to\infty}P(\psi^{q,s})=-\infty$.
@@
    For each $q>1$ the function $s\mapsto P(\psi^{q,s})$ is strictly increasing with $P(\psi^{q,0})\leqslant 0$ and $\lim_{s\to\infty}P(\psi^{q,s})=\infty$.
@@
    For each $q\neq 1$, there exists unique $s_{q}\in[0,\infty)$ so that $P(\psi^{q,s_{q}})=0$.
@@
    For each $q\in\mathbb{R}$ the function $s\mapsto P(\psi^{q,s})$ convex on connected components of $[0,\infty)\setminus\{1,2\}$.
@@
    For each $s\in[0,\infty)\setminus\{1,2\}$ the function $q\mapsto P(\psi^{q,s})$ convex on $\mathbb{R}$.
@@
Although the proof is a simple modification of [17, Lemma 2.1], we present the full details for the convenience of the reader. We prove the claims only for $s\in[0,2)$; the case $s\geqslant 2$ is left to the reader. Let $p,q\in\mathbb{R}$ and $s,t\in[0,2)$. Writing
@@
where $C\geqslant 1$ is the constant given by the quasi-Bernoulli assumption, we see that $0<\underline{\alpha}\leqslant\overline{\alpha}<1<K$. Furthermore, let
@@
and the function $(q,s)\mapsto P(\psi^{q,s})$ is thus clearly continuous on ×$\mathbb{R}\times[0,2)$. This shows (1). In particular, if $q<1$ and $s>t$, then the above estimate shows that $P(\psi^{q,s})-P(\psi^{q,t})\leqslant(s-t)(1-q)\log\overline{\alpha}(q,s,t)=(s-t)(1-q)\log\overline{\alpha}<0$, and if $q>1$ and $s>t$, we get $P(\psi^{q,s})-P(\psi^{q,t})\geqslant(s-t)(1-q)\log\overline{\alpha}>0$. These observations show (2)–(4). Notice that (5) follows immediately from (1).
Finally, let us prove (6) and (7). Fix $p,q\in\mathbb{R}$ and $10<\lambda<$ 1. Let $s,t\in[0,2)\setminus\{1\}$ be such that $\lceil s\rceil=\lceil t\rceil$. Since $\varphi^{\lambda t+(1-\lambda)s}(\mathtt{i})=\varphi^{t}(\mathtt{i})^{\lambda}\varphi^{s}(\mathtt{i})^{1-\lambda}$, we have
@@
for all $n\in\mathbb{N}$. The claims follow now by taking logarithm, dividing by $n$, and letting $n\to\infty$. ∎
Remark 3.2.
Although both coordinate functions $q\mapsto P(\psi^{q,s})$ and $s\mapsto P(\psi^{q,s})$ are convex on certain intervals by the assertions (6) and (7) of Lemma 3.1, the function $(q,s)\mapsto P(\psi^{q,s})$ is not necessarily convex. For example, if $10<\alpha_{2}<\alpha_{1}<$ 1 and
@@
In this case, the function $(q,s)\mapsto P(\psi^{q,s})$ is not convex since its Hessian is indefinite as it is an antidiagonal matrix having
@@
Let us next study the differentiability of the pressure. We will first recall some basic facts from convex analysis. Let $U\subset\mathbb{R}$ be an open set and let $f\colon U\to\mathbb{R}$ be convex. It is well known that such $f$ is continuous. We say that $G\in\mathbb{R}$ is a sub-derivative of $f$ at $x\in U$ if
@@
for all $y\in U$. It is straightforward to see that any sub-derivative is contained in $[f_{-}^{\prime}(x),f_{+}^{\prime}(x)]$, where $f_{-}^{\prime}(x)$ and $f_{+}^{\prime}(x)$ are the left and right derivatives of $f$ at $x$, respectively; see [20, Theorem 23.2]. Therefore, $f$ is differentiable at $x$ if and only if it has only one sub-derivative at $x$; see [20, Theorem 25.1].
@@
Let $(A_{1},\ldots,A_{N})\in GL_{2}(\mathbb{R})^{N}$ be dominated tuple of contractive matrices and $\mu\in\mathcal{M}(\Sigma)$ be a quasi-Bernoulli measure. If $(q_{0},s_{0})\in\mathbb{R}\times(0,\infty)\setminus\{1,2\}$ and $\nu$ is the equilibrium state for $\psi^{q_{0},s_{0}}$, then the partial derivatives of $(q,s)\mapsto P(\psi^{q,s})$ are
@@
for all $q\in\mathbb{R}$. Therefore, $-h(\mu,\nu)-\Lambda(\varphi^{s_{0}},\nu)$ is a sub-derivative of the convex function $q\mapsto Q(q,s_{0})$ at $q_{0}$. As the partial derivative $\partial_{q}Q(q,s_{0})|_{q=q_{0}}$ exists, we have $\partial_{q}Q(q,s_{0})|_{q=q_{0}}=-h(\mu,\nu)-\Lambda(\varphi^{s_{0}},\nu)$.
@@
for all all $s\in(0,2)\setminus\{1\}$ with $\lceil s\rceil=\lceil s_{0}\rceil$. Therefore, $(1-q_{0})\lambda_{\lceil s_{0}\rceil}(\nu)$ is a sub-derivative of the convex function $s\mapsto Q(q_{0},s)$ at $s_{0}$. As the partial derivative $\partial_{s}Q(q_{0},s)|_{s=s_{0}}$ exists, we have $\partial_{s}Q(q_{0},s)|_{s=s_{0}}=(1-q_{0})\lambda_{\lceil s_{0}\rceil}(\nu)$. ∎
Let us next show that $(q,s)\mapsto P(\psi^{q,s})$ is differentiable on ×$\mathbb{R}\times(0,\infty)\setminus\{1,2\}$ which then allows us to apply Proposition 3.3. To prove this requires tools from thermodynamic formalism. The following lemma allows us to employ Lemma 2.4 in this setting.

### v2

@@ 2.2 Lyapunov Dimension
We shall consider maps $\theta\colon\Sigma_{*}\to(0,\infty)$ which we refer to as potentials. We say that a potential $\theta$ is sub-multiplicative if $\theta(\mathtt{i}\mathtt{j})\leqslant\theta(\mathtt{i})\theta(\mathtt{j})$ for all $\mathtt{i},\mathtt{j}\in\Sigma_{*}$. A potential $\theta$ is super-multiplicative if the inverse $1/\theta$ is sub-multiplicative. We furthermore say that a potential $\theta$ is almost-multiplicative if there is a constant $C\geqslant 1$ such that $C\theta$ is sub-multiplicative and $C^{-1}\theta$ is super-multiplicative, and multiplicative if the constant $C$ can be chosen to be $1$. Let $\mathcal{M}_{\sigma}(\Sigma)$ denote the collection of all $\sigma$-invariant Borel probability measures on $\Sigma$. For a sub-multiplicative potential $\theta$ and $\nu\in\mathcal{M}_{\sigma}(\Sigma)$, we define
@@
The following well-known lemma guarantees that $\Lambda$ is well-defined.
@@
for $\nu$-almost all $\mathtt{i}\in\Sigma$. Furthermore, $\nu\mapsto\Lambda(\theta,\nu)$ defined on $\mathcal{M}_{\sigma}(\Sigma)$ is upper semi-continuous in the weak topology.
@@
The second claim follows from Kingman’s ergodic theorem; see e.g. [35, Theorem 10.1]. The third claim is a direct consequence of the first claim as each $\nu\mapsto\frac{1}{n}\sum_{\mathtt{i}\in\Sigma_{n}}\nu([\mathtt{i}])\log\theta(\mathtt{i})$ is continuous. ∎
Let $(A_{1},\ldots,A_{N})\in GL_{2}(\mathbb{R})^{N}$ and write $A_{\mathtt{i}}=A_{i_{1}}\cdots A_{i_{n}}$ for all $\mathtt{i}=i_{1}\cdots i_{n}\in\Sigma_{n}$ and $n\in\mathbb{N}$. By $A_{\mathtt{i}}^{-1}$ we mean $(A_{\mathtt{i}})^{-1}$. For $\mathtt{i}\in\Sigma_{*}$ we define $\alpha_{1}(\mathtt{i})$ and $\alpha_{2}(\mathtt{i})$ to be the lengths of the major and minor semi-axis of the ellipse $A_{\mathtt{i}}(D)$ respectively, where $D\subset\mathbb{R}^{2}$ is the unit disc. Note that $\alpha_{1}(\mathtt{i})=\|A_{\mathtt{i}}\|$ and $\alpha_{2}(\mathtt{i})=\|A_{\mathtt{i}}^{-1}\|^{-1}$ for all $\mathtt{i}\in\Sigma_{*}$, where $\|\cdot\|$ is the Euclidean operator norm. The potential $\mathtt{i}\mapsto\alpha_{1}(\mathtt{i})$ is thus sub-multiplicative and $\mathtt{i}\mapsto\alpha_{2}(\mathtt{i})$ is super-multiplicative. We define the Lyapunov exponents of $\nu\in\mathcal{M}_{\sigma}(\Sigma)$ by
@@
We say that a fully supported measure $\mu\in\mathcal{M}(\Sigma)$ is sub-multiplicative if the potential $\mathtt{i}\mapsto\mu([\mathtt{i}])$ is sub-multiplicative. The other definitions on potentials can be used with measures in a similar manner. Almost-multiplicative measures are more commonly known as quasi-Bernoulli measures, while multiplicative measures are called Bernoulli measures. The cross-entropy of a sub-multiplicative measure $\mu$ relative to $\nu\in\mathcal{M}_{\sigma}(\Sigma)$ is defined to be
@@
Since $x\log x\geqslant 0$ for all $x\in(0,1]$, we have $h(\nu)\geqslant 0$. To show the other inequality, fix $n\in\mathbb{N}$ and, as $\log x\leqslant x-1$ for all $x>0$, note that
@@ 2.3 Domination
We say that $\mathsf{A}=(A_{1},\ldots,A_{N})\in GL_{2}(\mathbb{R})^{N}$ is dominated if there exist constants $C>0$ and $0<\tau<1$ such that
@@
for all $\mathtt{i}\in\Sigma_{n}$ and $n\in\mathbb{N}$. Note that if $\mathsf{A}$ is dominated, then $\lambda_{2}(\nu)<\lambda_{1}(\nu)$ for all $\nu\in\mathcal{M}_{\sigma}(\Sigma)$. Let $\mathbb{RP}^{1}$ denote the real projective line, which is the set of straight lines going through the origin in $\mathbb{R}^{2}$ and which we identify with $0,\pi)$. We call a proper subset $\mathcal{C}\subset\mathbb{RP}^{1}$ a multicone if it is a finite union of closed projective intervals. We say that a multicone $\mathcal{C}\subset\mathbb{RP}^{1}$ is strongly invariant for $\mathsf{A}$ if $A_{i}\mathcal{C}\subset\mathcal{C}^{o}$ for all $i\in\{1,\ldots,N\}$, where $\mathcal{C}^{o}$ is the interior of $\mathcal{C}$. For example, the union of the first and third quadrants is strongly invariant for any tuple of positive matrices. Notice that if a multicone $\mathcal{C}\subset\mathbb{RP}^{1}$ is strongly invariant for $\mathsf{A}=(A_{1},\ldots,A_{N})$, then the closure of $\mathbb{RP}^{1}\setminus\mathcal{C}$ is a strongly invariant multicone for $(A_{1}^{-1}\ldots,A_{N}^{-1})$. By [13, Theorem B], $\mathsf{A}$ has strongly invariant multicone if and only if $\mathsf{A}$ is dominated. Therefore, if $\mathsf{A}$ is dominated, then the tuple $(A_{1}^{-1}\ldots,A_{N}^{-1})$ of inverses has a strongly invariant multicone. Furthermore, if $\mathsf{A}$ is dominated, then [14, Lemma 2.2] (see also [4, Corollary 2.4]) implies that the potential $\mathtt{i}\mapsto\alpha_{1}(\mathtt{i})$ is almost-multiplicative. Since $|\det(A_{\mathtt{i}})|=\alpha_{1}(\mathtt{i})\alpha_{2}(\mathtt{i})$ for all $\mathtt{i}\in\Sigma_{*}$ and the determinant is multiplicative, it follows that $\mathtt{i}\mapsto\alpha_{2}(\mathtt{i})=\alpha_{1}(\mathtt{i})^{-1}|\det(A_{\mathtt{i}})|$ is also almost-multiplicative.
@@
Let $\theta$ be an almost-multiplicative potential and $\nu\in\mathcal{M}_{\sigma}(\Sigma)$. If $\nu_{k}\to\nu$ in the weak topology, then
@@
Note that $X_{F}$ is uniformly perfect and hence has positive Hausdorff dimension unless it is a singleton. See [1, Lemma 2.3] for alternative characterizations of $X_{F}$. If $\nu\in\mathcal{M}_{\sigma}(\Sigma)$ is ergodic, then the measure $\nu_{F}=\Pi_{*}\nu$ on $X_{F}$ is called the Furstenberg measure with respect to $\nu$. Let $T\colon\Sigma\times\mathbb{RP}^{1}\to\Sigma\times\mathbb{RP}^{1}$ be the skew-product defined by
@@
If $\nu$ is Bernoulli, the measure $\nu\times\nu_{F}$ on $\Sigma\times\mathbb{RP}^{1}$ is $T$-invariant and ergodic. If $\nu$ is only quasi-Bernoulli, then $\nu\times\nu_{F}$ may not be invariant, but is equivalent to a $T$-invariant ergodic measure; see [6, Theorem 2.2] and [16, Lemma 3.1], respectively.
@@ 2.4 Equilibrium States
As in Lemma 2.1, the existence of the limit above and the equality are guaranteed by Fekete’s Lemma. By [22, Lemma 2.2], we see that
@@
is called the equilibrium state for $\theta$. If $\theta$ is an almost-multiplicative potential, then, by [29, Theorem 5] (or [12, Theorem 14]), there exists a unique equilibrium state for $\theta$ which furthermore is a fully supported quasi-Bernoulli measure.
@@
Let $(\theta_{k})_{k\in\mathbb{N}}$ be a sequence of sub-multiplicative potentials and let $\nu_{k}\in\mathcal{M}_{\sigma}(\Sigma)$ be an equilibrium state for $\theta_{k}$ for each $k\in\mathbb{N}$. If there exist a measure $\nu\in\mathcal{M}_{\sigma}(\Sigma)$ and a sub-multiplicative potential $\theta$ such that $\nu_{k}\to\nu$ in the weak topology and $\theta_{k}(\mathtt{i})^{1/|\mathtt{i}|}\to\theta(\mathtt{i})^{1/|\mathtt{i}|}$ uniformly in $\Sigma_{*}$ as $k\to\infty$, then
@@
Recall that, by [35, Theorem 8.2] and Lemma 2.1, $\limsup_{k\to\infty}h(\nu_{k})\leqslant h(\nu)$ and $\limsup_{k\to\infty}\Lambda(\theta,\nu_{k})\leqslant\Lambda(\theta,\nu)$. If $\varepsilon>0$, then the uniform convergence of $\theta_{k}$ implies that there exists $k_{0}\in\mathbb{N}$ such that $\log\theta(\mathtt{i})-\varepsilon|\mathtt{i}|\leqslant\log\theta_{k}(\mathtt{i})\leqslant\log\theta(\mathtt{i})+\varepsilon|\mathtt{i}|$ for all $\mathtt{i}\in\Sigma_{*}$ and
@@
for all $\mathtt{i}\in\Sigma_{*}$. Since $\alpha(\mathtt{i})\alpha_{2}(\mathtt{i})^{s-1}=\alpha_{1}(\mathtt{i})^{2-s}|\det(A_{\mathtt{i}})|^{s-1}$, the singular value function $\varphi^{s}$ is sub-multiplicative. Therefore, the pressure $P(\varphi^{s})$ is well-defined for all $s\geqslant 0$. By [23, Lemma 2.1], the function $s\mapsto P(\varphi^{s})$ defined on $0,\infty)$ is continuous and strictly decreasing, convex on intervals $[0,1]$, $[1,2]$, and $2,\infty)$, and there exists a unique $s\geqslant 0$ such that $P(\varphi^{s})=0$. This unique value is called the affinity dimension, denoted by $\operatorname{dim_{aff}}(\varphi^{s})$.
@@
By [24, Theorems 2.6 and 4.1], there exists an ergodic equilibrium state $\nu$ for $\varphi^{s}$. If $\mathsf{A}$ is dominated, then $\varphi^{s}$ is almost-multiplicative and $\nu$ is quasi-Bernoulli. It is easy to see that an equilibrium state $\nu$ has maximal possible Lyapunov dimension, $\operatorname{dim_{L}}(\nu)=\max\{\operatorname{dim_{L}}(\eta):\eta\in\mathcal{M}_{\sigma}(\Sigma)\}=\operatorname{dim_{aff}}(\varphi^{s})$.
@@ 3 Differentiability of the Pressure
is well-defined and there exists a unique ergodic equilibrium state for $\psi^{q,s}$, which is moreover a fully supported quasi-Bernoulli measure.
In this section, we prove that the function $(q,s)\mapsto P(\psi^{q,s})$ is differentiable on $\mathbb{R}\times(0,\infty)\setminus\{1,2\}$ and compute its partial derivatives. The differentiability of the pressure with respect to several real parameters in the diagonal case for almost-multiplicative potentials is addressed in [9]. We begin by collecting some elementary properties of the pressure function.
@@
    The function $(q,s)\mapsto P(\psi^{q,s})$ is continuous on $\mathbb{R}\times[0,\infty)$.
@@
    For each $q<1$, the function $s\mapsto P(\psi^{q,s})$ is strictly decreasing with $P(\psi^{q,0})\geqslant 0$ and $\lim_{s\to\infty}P(\psi^{q,s})=$-$\infty$.
@@
    For each $q>1$, the function $s\mapsto P(\psi^{q,s})$ is strictly increasing with $P(\psi^{q,0})\leqslant 0$ and $\lim_{s\to\infty}P(\psi^{q,s})=\infty$.
@@
    For each $q\neq 1$, there exists a unique $s_{q}\in[0,\infty)$ so that $P(\psi^{q,s_{q}})=0$.
@@
    For each $q\in\mathbb{R}$, the function $s\mapsto P(\psi^{q,s})$ is convex on connected components of $0,\infty)\setminus\{1,2\}$.
@@
    For each $s\in[0,\infty)\setminus\{1,2\}$, the function $q\mapsto P(\psi^{q,s})$ is convex on $\mathbb{R}$.
@@
Although the proof is a straightforward modification of [23, Lemma 2.1], we include the details for the convenience of the reader. We prove the claims only for $s\in[0,2)$; the case $s\geqslant 2$ is left to the reader. Let $p,q\in\mathbb{R}$ and $s,t\in[0,2)$. Write
@@
where $C\geqslant 1$ is the constant from the quasi-Bernoulli assumption. Then $0<\underline{\alpha}\leqslant\overline{\alpha}<1<K$. Furthermore, let
@@
and hence the function $(q,s)\mapsto P(\psi^{q,s})$ is continuous on $\mathbb{R}\times[0,2)$. This proves (1). In particular, if $q<1$ and $s>t$, then the above estimate shows that $P(\psi^{q,s})-P(\psi^{q,t})\leqslant(s-t)(1-q)\log\overline{\alpha}(q,s,t)=(s-t)(1-q)\log\overline{\alpha}<0$, and if $q>1$ and $s>t$, we get $P(\psi^{q,s})-P(\psi^{q,t})\geqslant(s-t)(1-q)\log\overline{\alpha}>0$. These observations give (2)–(4). Notice that (5) follows immediately from (1).
It remains to prove (6) and (7). Fix $p,q\in\mathbb{R}$ and $0<\lambda<1$. Let $s,t\in[0,2)\setminus\{1\}$ be such that $\lceil s\rceil=\lceil t\rceil$. Since $\varphi^{\lambda t+(1-\lambda)s}(\mathtt{i})=\varphi^{t}(\mathtt{i})^{\lambda}\varphi^{s}(\mathtt{i})^{1-\lambda}$, we have
@@
for all $n\in\mathbb{N}$. The claims follow by taking logarithms, dividing by $n$, and letting $n\to\infty$. ∎
Although the coordinate functions $q\mapsto P(\psi^{q,s})$ and $s\mapsto P(\psi^{q,s})$ are convex on certain intervals by Lemma 3.1(6) and (7), the function $(q,s)\mapsto P(\psi^{q,s})$ need not be convex. For example, if $0<\alpha_{2}<\alpha_{1}<1$ and
@@
In this case, the function $(q,s)\mapsto P(\psi^{q,s})$ is not convex because its Hessian is indefinite. Indeed, it is an antidiagonal matrix with
@@
Let us next study the differentiability of the pressure. We will first recall some basic facts from convex analysis. Let $U\subset\mathbb{R}$ be an open set and let $f\colon U\to\mathbb{R}$ be convex. It is well known that such a function is continuous. We say that $G\in\mathbb{R}$ is a sub-derivative of $f$ at $x\in U$ if
@@
for all $y\in U$. It is straightforward to see that any sub-derivative is contained in $f_{-}^{\prime}(x),f_{+}^{\prime}(x)]$, where $f_{-}^{\prime}(x)$ and $f_{+}^{\prime}(x)$ are the left and right derivatives of $f$ at $x$, respectively; see [33, Theorem 23.2]. Therefore, $f$ is differentiable at $x$ if and only if it has only one sub-derivative at $x$; see [33, Theorem 25.1].
@@
Let $(A_{1},\ldots,A_{N})\in GL_{2}(\mathbb{R})^{N}$ be a dominated tuple of contractive matrices and $\mu\in\mathcal{M}(\Sigma)$ be a quasi-Bernoulli measure. If $(q_{0},s_{0})\in\mathbb{R}\times(0,\infty)\setminus\{1,2\}$ and $\nu$ is the equilibrium state for $\psi^{q_{0},s_{0}}$, then the partial derivatives of $(q,s)\mapsto P(\psi^{q,s})$ are
@@
for all $q\in\mathbb{R}$. Therefore, $h(\mu,\nu)-\Lambda(\varphi^{s_{0}},\nu)$ is a sub-derivative of the convex function $q\mapsto Q(q,s_{0})$ at $q_{0}$. As the partial derivative $\partial_{q}Q(q,s_{0})|_{q=q_{0}}$ exists, we have $\partial_{q}Q(q,s_{0})|_{q=q_{0}}=-h(\mu,\nu)-\Lambda(\varphi^{s_{0}},\nu)$.
@@
for all $s\in(0,2)\setminus\{1\}$ with $\lceil s\rceil=\lceil s_{0}\rceil$. Therefore, $(1-q_{0})\lambda_{\lceil s_{0}\rceil}(\nu)$ is a sub-derivative of the convex function $s\mapsto Q(q_{0},s)$ at $s_{0}$. As the partial derivative $\partial_{s}Q(q_{0},s)|_{s=s_{0}}$ exists, we have $\partial_{s}Q(q_{0},s)|_{s=s_{0}}=(1-q_{0})\lambda_{\lceil s_{0}\rceil}(\nu)$. ∎
Let us next show that $(q,s)\mapsto P(\psi^{q,s})$ is differentiable on $\mathbb{R}\times(0,\infty)\setminus\{1,2\}$. This allows us to apply Proposition 3.2. The proof uses tools from thermodynamic formalism. The following lemma allows us to employ Lemma 2.4 in this setting.